Previous

Calendar

May

2008

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

27

JD118

28

JD119
Site Visit (JH): Details

Collected soil samples---soil starting to thaw out.

29

JD120

30

JD121

1

JD122

SUMMER JEEP RENTAL STARTS TODAY (start with green Jeep; gray Jeep is getting glass repaired)

2

JD123
CU classes end

green Jeep traded in for gray Jeep

3

JD124

gray Jeep leaves for Alamosa

4

JD125 (warm/pleasant)
Site Visit (SpB): Details , Photos (32)

at 10:50 MST, the support cable for the 25m REBs Q*7 sensor was loose; reconnected it (this might slightly affect the angle of this sensor).

from 11:30-12:30 MST, replaced 2-micron Nupro filter at LI-6262 inlet, also replaced nalgene at inlet

downloaded data from all loggers except UC and UCB

5

JD126

6

JD127

7

JD128

8

JD129
Ice/snow/rain, in-canopy csats appear to be iced up

9

JD130
Site Visit (JH): Details

Collected soil samples.

gray Jeep returns late from Alamosa, gets traded in for green Jeep (so gray Jeep's windows can get repaired)

note: road from MRS to C1 is now passable

10

JD131

11

JD132

12

JD133

13

JD134

14

JD135

15

JD136 (not too cold, heavy snow on and off)

Site Visit (SpB, JB): Details , Photos (28)

note: road from MRS to C1 was open, but snowy

fixed T3 on the tower

grabbed snow depths (see details)

started massive trailer cleaning

see details for cylinder removal/cal info

16

JD137

17

JD138

18

JD139

19

JD140 (warm)
Site Visit (JB, SO): Details

work on TRAM setup

20

JD141 (warm)
Site Visit (JB, SO): Details , Photos (5)

more work on TRAM setup

see details for snow depths

cleaned krypton optics

downloaded from soil, tcs, mrs logger and Noah's logger

N2 @ 350 psi, span @ 1000 psi

21

JD142
Site Visit (SpB, JB, SO, JH, NT, CB): Details , Photos (13)

more work on TRAM setup

22

JD143
Site Visit (JB, SO): Details

N2 at tower replaced (see details)

more work on TRAM setup

23

JD144
Site Visit (JB, SO, NT): Details

at 10:15 MST, Tower data system down (russter2 monitor not working)

at 14:55 MST, Tower data system back up

the monitor appears to be ok now.

more work on TRAM setup

see details for snow depths

24

JD145

25

JD146

26

JD147
CU Holiday (Memorial Day)
wet/overcast/rain.

27

JD148
Site Visit (NT, SL): Details

at 1:30 MST, power outage (data system goes down)

at 7:41 MST, power back up

at ~10:30 MST, quacker rebooted, data system back up

28

JD149

29

JD150

30

JD151

31

JD152